if you deposit $5,000 in an account paying 9% compounded daily, how much will you have in the account in 5 years? Compute the answer to the nearest cent.
Answer by rfer(16322)

A=P(1+.09/365)^1825
A=5000(1+.000246575)^1825
A=5000(1.568224214)
A=$7841.12